QUESTION 1

  1. Jordan Company produces basketballs and uses a standard costing system. Budgeted fixed overhead was $300,000. Rent changed during the year, causing actual fixed overhead to be $269,000. Jordan Company applies overhead on the basis of DLH. They projected 1,000,000 basketballs would be produced during the year. They actually produced 1,267,000 basketballs. The standard is 1 DLH/basketball. They actually used 1 DLH/basketball. What is the fixed overhead volume variance?

    (Please indicate overapplied fixed overhead with the letter "o" and underapplied fixed overhead with the letter "u")

QUESTION 2

  1. Druid Company makes a single product and uses a standard costing system that applies overhead on the basis of direct labor hours. They recently used 2,000 labor hours to produce 300 units. Their static budget indicated expectations of 500 units, 2,000 DLH and $60,000 in variable overhead. Actual VOH totaled $40,000. What was the variable overhead spending variance?

    Please indicate favorable variances with an "f" and unfavorable variances with a "u"

    (Round to the nearest dollar at the end of your calculations)

QUESTION 3

  1. Li Pong company uses a standard costing system. Last year they incurred $100,000 of Variable Overhead and $304,000 of Fixed Overhead and had the following variances before closing entries.

    FOH Budget Variance - $3,000 F

    FOH Volume Variance - $5,000 U

    VOH Spending Variance - $9,000 F

    VOH Efficiency Variance - $8,000 U

    How much overhead was applied to inventory over the course of the year?

    (Answer in dollars)
Answer #1

Solution 1:

Fixed overhead rate = $300,000 / 1000000 = $0.30 per DLH

Fixed overhead volume variance = Fixed overhead applied - Budgeted fixed overhead

= (1267000*$0.30) - $300,000 = $80,100 F

Solution 2:

Standard rate of variable overhead = $60,000 /2000 = $30 per DLH

Actual rate of variable overhead = $40,000 / 2000 = $20 per DLH

Variable overhead spending variance = (SR - AR) * AH = ($30 - $20) * 2000 = $20,000 F

Solution 3:

Overhead applied = Actual overhead + Favorable overhead variance - Unfavorable overhead variance

= ($100,000 + $304,000) + ($3,000 + $9,000) - ($5,000 + $8,000)

= $403,000